A pound of tin is worth $1 more than a pound of copper. Four pounds of tin are mixed with 6 pounds of copper to make bronze that sells for $3.65 per pound. How much is a pound of tin worth?
Answer by jorel1380(3719)

10x3.65=36.50
4(n)+6(n+1)=36.50
4n+6n+6=36.50
10n=30.50
n=3.05
n+1=4.05
A pound of tin is worth $4.05..